# DiC: Rethinking Conv3x3 Designs in Diffusion Models

### v1

@@ 1 Introduction
Unfortunately, diffusion models incorporating self-attention face significant computational overhead and latency when scaled up. The complex transformer architectures poses a considerable challenge for real-time, resource-constrained applications, where large-scale transformers are impractical due to their substantial time costs.
@@
In this work, we rethink the simple yet efficient fully convolutional diffusion model architecture that offers significant speed advantages. As a counterpart to self-attention, convolution is widely regarded as hardware-friendly, with excellent support across various platforms. Among the many convolutional operations, the canonical 3x3 stride-1 convolution stands out for its exceptional speed, largely due to widespread hardware optimization techniques such as Winograd acceleration [22]. For instance, works like RepVGG [11] have demonstrated how simple convolutions can achieve both high efficiency and competitive performance.
@@
To address this, we propose a series of adaptations to conventional ConvNets, forming a roadmap (shown in Fig. 1) that tailors them specifically for a scaled-up diffusion model. Firstly, we focus on refining the model architecture to better leverage Conv3x3. Existing designs fall into three categories: isotropic architectures like DiT [26], isotropic architectures with skips like U-ViT [1], and encoder-decoder hourglass structures used in canonical CNN & self-attention hybrid models. Empirical observations reveal that the hourglass architecture proves more effective for pure Conv3x3 models, because the downsampling and upsampling in encoders expand the receptive field that makes up for the narrow scope of Conv3x3. Skip connections are also observed crucial, but traditional blockwise skips for cheap convolution block scale-ups are too dense to be efficient for large ConvNets. To address this, we introduce sparse skip connections that reduces the number of skips while ensuring essential information flows efficiently from encoder to decoder.
Besides, we focus on improving conditioning to fit the hourglass ConvNet better. In existing models, a single set of condition embeddings is often mapped to different blocks across various stages. However, in encoder-decoder ConvNets, each stage operates in distinct feature spaces. Hence, we introduce stage-specific embeddings, ensuring each stage uses independent, non-overlapping condition embedding tables. We also carefully inspect the position for embedding injection, and we suggest inject conditions mid-block for better performance. Additionally, we borrow conditional gating on feature-maps. Further, we refined the internal block structure replacing all activation functions with GELU.
With these enhancements, we propose DiC, a diffusion model made up of Conv3x3. The architectural and conditioning improvements collectively enable our model to achieve outstanding results while maintaining its speed advantage. Experiments on various scales and settings demonstrate DiC’s performance advantage over Diffusion Transformers as well as its high throughput characteristics.
@@ 2 Related Work
Self-Attention in Diffusion Architectures. Self-attention has become a cornerstone of modern diffusion architectures, thanks to its remarkable effectiveness in capturing long-range dependencies. Early UNet-based diffusion models [18, 31] incorporated self-attention layers in the higher stages of the network, where capturing global context is most critical. Dhariwal et al. took this further with ADM [10], extending self-attention across all stages, which significantly boosted the model’s generative capabilities.
More recent works, such as DiT [26] and U-ViT [1], have fully embraced transformer-based architectures, replacing traditional U-Net backbones with full self-attention designs. These models not only achieve superior performance but also exhibit excellent scalability, particularly at large model size and FLOPs. Most following works like PixArt-$\alpha$ [2] and Stable Diffusion 3 adopt the full-transformer backbone design that demonstrates outstanding Text-to-Image generation capabilities.
Beyond these, some works continue to explore and refine the use of self-attention and transformer architectures in diffusion models, further pushing the boundaries of efficiency in transformer-based generative modeling. For instance, [30, 3, 33] looks at downsampled tokens for self-attention; [37] uses linear attention. Other works [38, 32] opt to State Space Models for diffusion, but SSMs suffer in terms of latency for ordinary-shaped images [32].
In contrast to these works, we push architectural simplicity to the extreme by designing an entire diffusion model using 3x3 convolutions. With carefully crafted structural optimizations, this streamlined network not only capitalizes on the speed advantages of 3x3 convolutions but also achieves diffusion generation performance on par with transformer-based models.
@@ 3.4 Minor Modification
We re-visit the activation function across the entire model. Inspired by ConvNeXt [24], we replaced the commonly used SiLU activation with GELU, which is a standard choice in transformer architectures. On top of other modifications, this change contributed to a modest but consistent improvement in performance.
@@ 4 Experiments
In this section, we provide experiments on various settings to demonstrate the outstanding performance of the proposed DiC model.
@@ 4.1 Experiment Setups
Model Configs. To benchmark against existing DiT [26] models, we designed our architectures at different scales: Small (S), Big (B), and Extra Large (XL), which align closely with DiT-S/2, DiT-B/2, and DiT-XL/2 in terms of both FLOPs and parameter count, as shown in Table 3. Additionally, we introduce a larger model, Huge (H), for further scalability testing. With 32.8M parameters and 5.9G FLOPs, DiC-S provides a lightweight solution, similar to DiT-S/2, with an efficient [6,6,5,6,6] encoder-decoder configuration. DiC-B scales up to 129.5M parameters and 23.5G FLOPs, which is comparable to DiT-B/2, leveraging a wider channel size of 192. DiC-XL has 702.3M parameters and 116.1G FLOPs, whose configuration rivals DiT-XL/2, with an enhanced encoder-decoder depth of [7,7,8,7,7]. DiC-H is further scaled up to [14,14,10,14,14] and is designed to evacuate the potentials of DiC models.
@@ 4.2 Evaluating DiC on the Standard DiT Setting
Table 4: Comparing DiCs against DiTs across various model sizes. We compare models trained for 400K iterations with the standard training hyperparameters of DiT. DiCs at all scales outperform DiTs at considerable margins.
@@
Figure 4: Comparing FID performance and FLOPs between the class of DiT and DiC models. Marker sizes represents relative model sizes.
@@ 4.3 DiC on Larger Images
Table 7: Comparing DiCs against DiTs on ImageNet 512$\times$512 generation. We compare models trained for 400K iterations with the standard training hyperparameters of DiT.
@@ 4.4 Scaling to Excellence
In this section, we look at the scaled-up model size with increased training iterations to discover the potential of DiC models.
ImageNet 256$\times$256, Scale Up
@@
Table 8: Fast Convergence of DiC. The unconditional generation performance of both models on ImageNet 256$\times$256 consistently improves as training progresses.
Unconditional Generation. Table 8 presents the performance of DiC-H on ImageNet 256$\times$256 as training progresses. The performance of DiC-H improves steadily as the training progresses. At 600K steps, FID decreases to 9.73, matching the performance of DiT-XL/2 at 7M training steps; at 800K steps, DiC-H shows the best performance with an FID of 8.96. This table demonstrates the fast convergence of DiC-H, where both FID and IS consistently improve with training.
ImageNet 256$\times$256, Scale Up, Conditional
@@
Figure 5: Samples generated by DiC at 2M iterations. The samples are generated following the setting of DiT, at $cfg=4$. Best viewed on screen.
Conditional Generation. Table 9 presents the comparison of various models’ performance on ImageNet 256$\times$256 as training iteration is scaled up. The key metrics presented are Throughput (TP), Batch Size $\times$ Iterations (BS×$\times$×Iter), and FID (Fréchet Inception Distance), a commonly used measure of generative model quality.
The baseline model DiT-XL/2 achieves a throughput of 66.8, processes a batch size of 256 for 7 million iterations, and yields an FID score of 2.27. The U-ViT-H model, while slightly lower in throughput at 63.9, uses a larger batch size of 1024, resulting in a slightly higher FID of 2.29. In comparison, our proposed DiC-H model shows a marked improvement when trained with a modest setting of batch size of 256 over 2 million iterations. DiC-H could achieve an FID of 2.25, with a throughput of 160.8. These results reveal that CNNs could reach performance that are no worse than Diffusion Transformers.
@@ 5 Conclusion
In this work, we rethink the design of Conv3x3 in the context of diffusion models and introduced DiC, a purely 3x3-convolutional diffusion model. We initially explored various state-of-the-art diffusion architectures and found that the Encoder-Decoder U-Net structure performed the best. Building on this, we proposed architectural improvements, such as the sparse skip connection, to further enhance performance.
In addition to architecture, a well-designed conditioning mechanism is essential. We used independent conditional embedding tables for each stage, allowing for more flexible and efficient conditioning. Furthermore, we adjusted the position of the condition injection, moving it from the beginning of the block to the middle, which showed improved results. Lastly, we incorporated conditional gating from DiT to refine the conditioning process. We also made adjustments to the activation function, optimizing it for better performance. Based on these improvements, our proposed Conv3x3 diffusion model not only achieves superior generative performance compared to Diffusion Transformers but also does so with significantly higher throughput.
@@ 6 Additional Experiments
More Traing Iterations on ImageNet 512x512. We have extended the training iterations for some limited number of iterations, shown in Tab. 11. Both DiC-XL and DiC-H could outperform DiT-XL/2 at much fewer training iterations while maintaining a speed advantage.
ImageNet 512$\times$512, Scale Up

### v2

@@ 1 Introduction
Unfortunately, diffusion models incorporating self-attention face significant computational overhead and latency when scaled up. Complex transformer architectures pose a considerable challenge for real-time, resource-constrained applications, where large-scale transformers are impractical due to their substantial time costs.
@@
In this work, we rethink the simple yet efficient fully convolutional diffusion model architecture that offers significant speed advantages. As a counterpart to self-attention, convolution is widely regarded as hardware-friendly, with excellent support across various platforms. Among the many convolutional operations, the canonical 3x3 stride-1 convolution stands out for its exceptional speed, largely due to widespread hardware optimization techniques such as Winograd acceleration [26]. For instance, work like RepVGG [12] have demonstrated how simple convolutions can achieve both high efficiency and competitive performance.
@@
To address this, we propose a series of adaptations to conventional ConvNets, forming a roadmap (shown in Fig. 1) that tailors them specifically for a scaled-up diffusion model. Firstly, we focus on refining the model architecture to better leverage Conv3x3. Existing designs fall into three categories: isotropic architectures like DiT [30], isotropic architectures with skips like U-ViT [1], and encoder-decoder hourglass structures used in canonical CNN & self-attention hybrid models. Empirical observations reveal that the hourglass architecture is more effective for pure Conv3x3 models, because downsampling and upsampling in encoders expand the receptive field that makes up for the narrow scope of Conv3x3. Skip connections are also observed crucial, but traditional blockwise skips for cheap convolution block scale-ups are too dense to be efficient for large ConvNets. To address this, we introduce sparse skip connections that reduces the number of skips while ensuring essential information flows efficiently from encoder to decoder.
Besides, we focus on improving conditioning to fit the hourglass ConvNet better. In existing models, a single set of condition embeddings is often mapped to different blocks across various stages. However, in encoder-decoder hourglass ConvNets, each stage operates in distinct feature spaces. Hence, we introduce stage-specific embeddings, ensuring each stage uses independent, non-overlapping condition embedding tables. We also carefully inspect the position for embedding injection, and we suggest inject conditions mid-block for better performance. Additionally, we borrow conditional gating on feature-maps. In addition, we refined the internal block structure, replacing all activation functions with GELU.
With these enhancements, we propose DiC, a diffusion model made up of Conv3x3. The architectural and conditioning improvements collectively enable our model to achieve outstanding results while maintaining its speed advantage. Experiments on various scales and settings demonstrate DiC’s performance advantage over Diffusion Transformers as well as its high-throughput characteristics.
@@ 2 Related Work
Self-Attention in Diffusion Architectures. Self-attention has become a cornerstone of modern diffusion architectures, thanks to its remarkable effectiveness in capturing long-range dependencies. Early UNet-based diffusion models [19, 36] incorporated Self-Attention layers in the higher stages of the network, where capturing global context is most critical. Dhariwal et al. took this further with ADM [11], extending self-attention across all stages, which significantly increased the model’s generative capabilities.
More recent works [1, 30, 21] have fully embraced transformer-based architectures, replacing traditional U-Net backbones with full self-attention designs. These models not only achieve superior performance, but also exhibit excellent scalability, particularly at large model sizes and FLOPs. Most following work like PixArt-$\alpha$ [2] and Stable Diffusion 3 adopt the full transformer backbone design that demonstrates outstanding Text-to-Image generation capabilities. Beyond these, some works continue to explore and refine the use of self-attention and transformer architectures in diffusion models, further pushing the boundaries of efficiency in transformer-based generative modeling. For instance, [35, 3, 38] looks at downsampled tokens for self-attention; [43] uses linear attention. Other works [45, 37] opt for State Space Models for diffusion, but SSMs suffer in terms of latency for ordinary-shaped images [37]. Some other work [20, 6, 22, 38] looks at a combination of U-Net and Self-Attention. In contrast to these works, we push architectural simplicity to the extreme by designing an entire diffusion model using 3x3 convolutions. With carefully crafted structural optimizations, this streamlined network not only capitalizes on the speed advantages of 3x3 convolutions but also achieves diffusion generation performance on par with transformer-based models.
@@ 3.4 Minor Modification
We re-visit the activation function across the entire model. Inspired by ConvNeXt [28], we replaced the commonly used SiLU activation with GELU, which is a standard choice in transformer architectures. On top of other modifications, this change contributed to a modest but consistent improvement in performance. We are aware of potentially better activations [31, 44], but we stick to GeLU for simplicity.
@@ 4 Experiments
In this section, we provide experiments on various settings to demonstrate the outstanding performance of DiC model.
@@ 4.1 Experiment Setups
Model Configs. To benchmark against existing DiT [30] models, we designed our architectures at different scales: Small (S), Big (B), and Extra Large (XL), which align closely with DiT-S/2, DiT-B/2, and DiT-XL/2 in terms of both FLOPs and parameter count, as shown in Table 3. Additionally, we introduce a larger model, Huge (H), for further scalability testing. With 32.8M parameters and 5.9G FLOPs, DiC-S provides a lightweight solution, similar to DiT-S/2, with an efficient [6,6,5,6,6] encoder-decoder configuration. DiC-B scales up to 129.5M parameters and 23.5G FLOPs, which is comparable to DiT-B/2, leveraging a wider channel size of 192. DiC-XL has 702.3M parameters and 116.1G FLOPs, whose configuration rivals DiT-XL/2, with an enhanced encoder-decoder depth of [7,7,8,7,7]. DiC-H is further scaled up to [14,14,10,14,14] and is designed to evaluate the potentials of DiC models.
@@ 4.2 Evaluating DiC on the Standard DiT Setting
Table 4: Comparing DiCs against DiTs across various model sizes. We compare models trained for 400K iterations with the standard setting of DiT. DiCs at all scales outperform DiTs at considerable margins.
@@
ImageNet 256$\times$256, 400K
[…]
DiT-XL/2 1.5 118.6 6.24 150.10
U-ViT-XL 1.5 113.0 5.66 170.62
DiC-XL 1.5 116.1 (57.2) 3.89 224.20
Figure 4: Comparing FID performance and FLOPs between DiTs and DiCs. Marker sizes represents relative model sizes.
@@ 4.3 DiC on Larger Images
Table 7: Comparing DiCs against DiTs on ImageNet 512$\times$512 generation. We compare models trained for 400K iterations with the standard setting of DiT. Please refer to the appendix for results from longer training iterations.
@@ 4.4 Scaling to Excellence
In this section, we look at the scaled-up model (DiC-H) with increased training iterations, and the combination with diffusion techniques to discover the potential of DiC.
ImageNet 256$\times$256, Scale Up, w/o cfg
@@
Table 8: Fast Convergence of DiC. The generation performance without classifier guidance of both models on ImageNet 256$\times$256 consistently improves as training progresses.
Under DiT’s plainest no-CFG setting. Table 8 presents the performance of DiC-H on ImageNet 256$\times$256 as training progresses. The performance of DiC-H improves steadily as the training progresses. At 600K steps, FID decreases to 9.73, matching the performance of DiT-XL/2 at 7M training steps; at 800K steps, DiC-H shows the best performance with an FID of 8.96. This table demonstrates the fast convergence of DiC-H, where both FID and IS consistently improve with training.
ImageNet 256$\times$256, Scale Up, w/ cfg
@@
Figure 5: Samples generated by DiC at 2M iterations. The samples are generated following the setting of DiT, at $cfg=4$.
Towards excellence with classifier-free guidance. Table 9 presents the comparison of various models’ performance on ImageNet 256$\times$256 when training iteration is scaled up and classifier-free guidance is applied.
The baseline model DiT-XL/2 achieves a throughput of 66.8, processes a batch size of 256 for 7 million iterations, and yields an FID score of 2.27. The U-ViT-H model, while slightly lower in throughput at 63.9, uses a larger batch size of 1024, resulting in a slightly higher FID of 2.29. In comparison, our proposed DiC-H model shows a marked improvement when trained with a modest setting of batch size of 256 over 2 million iterations. DiC-H could achieve an FID of 2.25, with a throughput of 160.8. These results reveal that CNNs could reach performance that is no worse than Diffusion Transformers, and they have a great throughput advantage. Building upon this SiT flow-based framework, we also combine DiC with the State-of-the-Art technique of Representation Alignment (REPA) [46, 39] to achieve much faster convergence. The details are in the appendix.
@@ 5 Conclusion
In this work, we rethink the design of Conv3x3 in the context of diffusion models and introduced DiC, a purely 3x3-convolutional diffusion model. We initially explore various state-of-the-art diffusion architectures and find that the Encoder-Decoder U-Net structure perform the best. Building on this, we propose the sparse skip connection to reduce redundant block-wise connections and enhance efficiency of U-Net. In addition to architecture, we redesign the conditioning mechanism. We introduce stage-specific independent embeddings where each embedding is responsible for addressing conditioning within its assigned stage. We also propose mid-block condition injection and conditional gating, which are beneficial from empirical obvervations. Based on these improvements, our proposed Conv3x3 diffusion model not only achieves superior generative performance compared to Diffusion Transformers but also does so with significantly higher throughput.
@@
Acknowledgement. This work is supported by the National Key R&D Program of China under grant No. 2022ZD0160300 and the National Natural Science Foundation of China under grant No. 62276007. We gratefully acknowledge the support of MindSpore, CANN, and Ascend AI Processor used for this research.
@@ 6 Additional Experiments
More Traing Iterations on ImageNet 512$\times$512. We have extended the training iterations for some limited number of iterations, shown in Tab. 11. Both DiC-XL and DiC-H could outperform DiT-XL/2 at much fewer training iterations while maintaining a speed advantage. As is shown in Tab. 12, DiC models could also perform better than DiT-XL at 3M iterations, besides the complexity and speed advantage of Conv3x3 especially on larger images.
ImageNet 512$\times$512, Scale Up, w/o cfg
@@
ImageNet 512$\times$512, 3M, cfg=1.5
[…]
DiT-XL/2 524.7 16.2 3.04 240.82
DiC-XL 464.3 (228.7) 84.2 3.04 271.77
DiC-H 817.2 (388.4) 53.3 2.96 293.54
Table 12: The performance of DiC on larger images at 3M iterations. DiC models could achieve better performance and higher throughput than DiT-XL on ImageNet 512$\times$512.
@@
Scaling plots. We visualize the scaling curve of DiC-XL and DiC-H as shown in Fig. 7.
Figure 7: The scaling curve of DiC training. The FID scores are recorded once every 200K iterations in the first 1M training iterations. The scaling effect of DiC as model gets larger is obvious from the plot.
Comparing with more architectures. Apart from the architectures mentioned in Tab. 4, we are aware of some other competitive architectures including Simple Diffusion [20], RIN [21], EDM2 [22], and HDiT [6]. However, we find difficulty in comparing these methods with DiC: DiC is mainly focused on 256-sized latent diffusion, following DiT [30] and SiT [29]; these work, on the other hand, focuses on large (mostly pixel-space) diffusion; and they require large training costs to reach SOTA FIDs (e.g. EDM2 requires the training cost of 939.5-2147.5M img, which is around 4M to 8M iterations in our setting). We try to align HDiT, RIN, and EDM2 to our setting (under the training framework of DiT; in order to keep FLOPs aligned with DiC for fair comparison, we increase the depths and widths of these models). Results turns out that these methods either converges slowly (RIN, EDM2) or completely fails (HDiT).
Details regarding Representation Alignment on DiC. We consider applying Representation Alignment (REPA) [46] (using its variant U-REPA [39] tailored for U-Net) to achieve faster convergence. We use the standard training hyperparameter for REPA. For sampling, we use $cfg=1.8$ for SDE, and $cfg=2$ for ODE, both equipped with guidance interval (following the default setting of the official codebase). Amazingly, DiC-XL could reach an FID of 1.74 after 1M training steps with the help of REPA, as shown in Tab. 13.
ImageNet 256$\times$256, REPA
[…]
DiC-XL+U-REPA 1M ODE 1.74
DiC-XL+U-REPA 1M SDE 1.75
Table 13: Fast Convergence of DiC with the help of REPA. DiC-XL could achieve 1.75 FID after 1M training iterations.